OpenStudy (anonymous):

how to enter the equation of the circle in standard form with center and radius given. Center (8, 0), radius =

OpenStudy (anonymous):

radius=\[\sqrt{3}\]

OpenStudy (lgbasallote):

equation of circle is \[(x-h)^2 + (y-k)^2 = r^2\] right? so just substitute

OpenStudy (anonymous):

similar (almost identical) to the last one

OpenStudy (lgbasallote):

maybe he's confused about the radical?